Summary Title: Pushing the limits
Author’s name: Andrew Wheeler
Mission Status: Nominal
Sol Activity Summary: Another clear and cold (-4C) morning. Perfect for the solar observatory and EVAs. A morning EVA to Pooh’s Corner allowed a team to revisit a site last observed 18 years ago to gauge arid land degradation with a view to develop a geoheritage strategy. Concurrent with this, more robot rover traction trials were conducted over the different terrain. This EVA was most successful. The afternoon EVA sought to push the rovers to the maximum of their endurance under the current conditions, in an attempt to reach White Moon at the junction of Cow Dung Road and Brahe Highway. Here, gypsum crystals are exposed at the surface and these provide excellent targets for spectral analysis. Under the conditions, the rovers made it to approximately 800m short of the target. This distance was easily walked and 20 minutes later the team were there. Spectra were taken, samples collected and, in too short a time, the team headed back to the MDRS. However, the journey was interrupted by the appearance of a pair of (juvenile?) pronghorn antelope who crossed our path. What are they doing on Mars?. During all of this, the sun was continuously imaged in the observatory and an enhanced composite produced. Cheesy garlic scones for lunch and jambalaya for dinner was greatly enjoyed by all.
Look Ahead Plan: A day of consolidation. With no EVAs planned (except for a short one to redeploy the solar logger), we will use the time to crunch data, process samples and collect laboratory spectra for comparison to field data.
Anomalies in work: Nothing to report.
Weather: Clear morning at -4 Celsius rising to mid teens during the day. No clouds and no precipitation. Humidity continuously dropping from near 70% at sunrise to 30% in the afternoon. The wind was faint to gentle during the day.
Crew Physical Status: Nothing to report.
EVA: EVA 15 with Clare, Steve and Alex in search of anhydrite yardangs near Pooh’s Corner complimented with robot rover trials.
EVA 17 with Andrew, Steve and Scott to White Moon to push the rovers to their limits and collect spectra and samples for analyses.
Reports to be filed: Sol Summary, Journalist’s Report, Greenhab Report, Operations Report (including an HSO report), EVA report, EVA Request(s), Astronomer’s report.
Support Requested: Nothing to request.